OpenStudy (anonymous):

Given and , find the length of QS and TV. A. 2 B. 9 C. 8 D. 20 i need help really bad ill give a medal

OpenStudy (jdoe0001):

can you post a picture of the material? just a quick screenshot will do

OpenStudy (anonymous):

QS and TV both equal 8

OpenStudy (jdoe0001):

\(\bf \triangle QRS \cong \triangle TUV\) implies that both triangles are pretty much equal, not just similar that means, all their sides are equal notice, QS on the first triangle is correspondent to TV on the second triangle thus 

OpenStudy (anonymous):

so they are both 8?

OpenStudy (jdoe0001):

so, if we solve for "v" \(\bf \triangle QRS \cong \triangle TUV \\ \quad \\ 3v+2=7v-6\implies 2+6=7v-3v\implies 8=4v \\ \quad \\ \cfrac{\cancel{8}}{\cancel{4}}=v\implies ?=v\)

OpenStudy (jdoe0001):

so.... what does that give you for "v"?

OpenStudy (anonymous):

yes they are both equal to 8

OpenStudy (jdoe0001):

k